# Bridge the Modality and Capacity Gaps in Vision-Language Model Selection

### v1

@@ 2.3 Analysis of the Two Gaps in LOVM
There are two main challenges that limit the application of the aforementioned paradigms in LOVM. The first one is the Modality Gap across different modalities’ features in VLM, and the second is the Capacity Gap between VLM’s average performance and dataset-specific performance.
@@
Modality Gap. As described in subsection 2.2, some methods like H-Score, LogME, and ModelGPT utilize the ChatGPT generated auxiliary texts $\tilde{D}_{\mathcal{T}}$ as image proxies to calculate metrics that measure the zero-shot accuracy on the target dataset $\mathcal{T}$. In other words, the zero-shot classification ability across text and image modalities is estimated by the intra-modality classification ability. The latent assumption is that the generated texts and their corresponding images are closely aligned in VLM’s feature space. However, this assumption is difficult to meet (Liang et al., 2022), and VLM’s features are more likely to cluster according to their modalities. In particular, we define the modality gap vector $\bm{g}$ between the features of an image-text pair $({\bm{x}}_{i},t_{i})$ as:
$\bm{g}_{m,i}\coloneqq f_{m}^{I}({\bm{x}}_{i})-f_{m}^{T}(t_{i})\;.$ (9)
Values in the gap vector are far from zero in most cases. We name this phenomenon as Modality Gap in LOVM, which makes the scores on $\tilde{D}_{\mathcal{T}}$ hard to reveal the true zero-shot image classification capability of a VLM on a given dataset.
@@
We conduct a validation experiment on ImageNet with 35 VLMs provided by (Zohar et al., 2023). For each VLM, we first generate 50 auxiliary texts per class as $\tilde{D}_{\mathcal{T}}$ and then calculate the predicted Top-1 accuracy via Equation 8. Next, we use test images to calculate the VLM’s true Top-1 accuracy. The consistency between the predicted Top-1 accuracy and true zero-shot image classification accuracy $p_{m,\mathcal{T}}$ is measured by the Kendall Rank Correlation ($\tau$, higher is better) and Mean Absolute Error (MAE, lower is better). The results are shown in Figure 2, where each point represents a VLM. It can be observed that the predicted accuracy derived from auxiliary texts $\tilde{D}_{\mathcal{T}}$ does not closely match the true zero-shot accuracy, indicating that these generated auxiliary texts in $\tilde{D}_{\mathcal{T}}$ are not effective proxies for images.
@@
To make the auxiliary texts act as better image proxies, one intuitive idea is to estimate the gap vector $\bm{g}$ for each image-text pair. Given the vector, we can add it to the feature $f_{m}^{T}(t_{i})$ of the text $t_{i}$ to eliminate the modality gap, which may further lead to more accurate scores $s_{m,\mathcal{T}}^{i}$ in Equation 7. However, the gap vector cannot be calculated directly without the images from the target dataset. Furthermore, gap vectors for different classes are diverse, so using a shared vector across all datasets may not be a good choice in LOVM.
Figure 3: The workflow of Swab. Swab first uses optimal transport to construct the transport matrix $\bm{\gamma}^{*}\in\mathbb{R}^{k_{\mathcal{S}}\times k_{\mathcal{T}}}$ based on the textual semantic similarity between the open-source datasets’ classes $C_{\mathcal{S}}=\{c_{1}^{\mathcal{S}},\cdots,c_{k_{\mathcal{S}}}^{\mathcal{S}}\}$ and target dataset’s classes $C_{\mathcal{T}}=\{c_{1}^{\mathcal{T}},\cdots,c_{k_{\mathcal{T}}}^{\mathcal{T}}\}$. Then Swab uses the transport matrix to estimate the VLM’s class-specific gap vectors $\{\bm{g}_{m,1},\cdots\}$ on the target dataset $\mathcal{T}$ based on VLM’s gap vectors $\bm{G}_{m}^{\mathcal{S}}\in\mathbb{R}^{k_{\mathcal{S}}\times d}$ on the open-source datasets. The estimated gap vectors are used to modify text samples for acting as better image proxies, which are input to the Ranker Model $f_{R}$ to predict the $m$-th VLM’s performance $\hat{p}_{m,\mathcal{T}}^{1}$ on the target dataset. Besides, Swab also predicts VLM’s performance based on the VLM’s class-specific ranking $\bm{r}_{m}^{\mathcal{S}}\in\mathbb{R}^{k_{\mathcal{S}}}$ on open-source datasets using the transport matrix $\bm{\gamma}^{*}$. At last, Swab ensembles these predicted performance $\hat{p}_{m,\mathcal{T}}^{1}$ and $\hat{p}_{m,\mathcal{T}}^{2}$ to get the VLM’s final ranking prediction.
@@
Capacity Gap. To select one VLM from the model zoo given a target dataset, one direct approach is to select the “strongest” VLM in all cases. For example, we may first estimate the VLM’s zero-shot classification ability on open-source datasets and then utilize the VLM with the highest performance, which is described in subsection 2.2. Will a VLM’s average ranking on the open-source datasets reveal its true ranking on the target dataset? Our empirical analyses indicate that there exists a discrepancy between the VLM’s overall ranking and its ranking on a specific target dataset. We name the difference between the average ability and the specific ability as the Capacity Gap, which results from the fact that a VLM’s performance fluctuates significantly across various datasets.
@@
To verify the claim, we test 35 VLMs on 23 target datasets provided by (Zohar et al., 2023) and obtain the rankings of each VLM across these datasets. Based on these ranking results, we calculate the average standard deviation and the mean difference between each VLM’s maximum and minimum ranking. The details are illustrated in the right part of Figure 2. The high value of average standard deviation and range difference demonstrate that a VLM’s zero-shot classification ability depends on the property of the target dataset. For example, the mean difference between the maximum and the minimum ranking of one VLM is 28, while the total number of evaluated VLMs is 35. Therefore, the top-performing VLM in one dataset could likely be among the worst in another.
@@
To bridge such a capacity gap, one possible solution is to take account of the ranking of all VLMs on a related dataset. In other words, instead of using the general performance ranking, the specific ranking on some sampled datasets from open-source datasets may be helpful. The main challenge is to figure out which open-source dataset is similar to the target one, and transform the VLM’s ranking on that dataset to the target dataset.
@@
Summary. We emphasize two kinds of gaps in LOVM, i.e., the modality gap across features of different modalities generated by a VLM, and the capability gap between a VLM’s overall ranking and its ranking given a specific target dataset. Both two gaps pose obstacles to some methods, such as LogME and ModelGPT, and degrade their abilities in VLM selection. Moreover, those intuitive approaches to bridge the gaps still face challenges.
@@ 3 VLM Selection with Gap Bridging
To mitigate the impact of both gaps on LOVM and integrate non-learning-based and learning-based LOVM methods, we propose VLM Selection With gAp Bridging (Swab). The key idea of Swab lies in bridging the modality and capacity gaps by leveraging useful class-level statistics of VLMs based on open-source datasets. By measuring the textual similarity between the target dataset’s class names and those in open-source datasets, we construct a bridge matrix, which indicates how classes from the open-source datasets can depict a particular VLM’s characteristics for the target dataset. We then estimate the gap vectors between image and text modalities, which rectifies the text-derived scores in ModelGPT. In addition, we assess the VLM’s performance ranking for a specified dataset. Two estimations are merged to achieve a more accurate language-only VLM selection. The workflow of Swab is illustrated in Figure 3.
3.1 Construct the Bridge Using Optimal Transport
@@
Benefiting from the open-source datasets, some useful class-level statistics, such as modality gap vectors and zero-shot classification accuracy of a certain VLM, could be calculated, which may help the ability estimation of a VLM on the target dataset. Instead of reusing the class-level statistics in a uniform manner, our Swab automatically determines the semantic relationship between open-source datasets’ classes and target dataset’s classes with Optimal Transport (Cuturi, 2013; Peyré & Cuturi, 2019).
@@
Recall that the sets of class names of the open-source datasets and the target dataset are $C_{\mathcal{S}}=\{c^{\mathcal{S}}_{i}\}^{k_{\mathcal{S}}}_{i=1}$ and $C_{\mathcal{T}}=\{c^{\mathcal{T}}_{i}\}^{k_{\mathcal{T}}}_{i=1}$, respectively. The semantic relationship between two classes could be measured by the textual similarity between their class names. In detail, we use a pre-trained text encoder $\phi$ (e.g., MPNet (Song et al., 2020)), which extracts same-dimensional features for all previous class names, i.e., $\{\phi(c_{1}^{\mathcal{S}}),\cdots,\phi(c_{k_{\mathcal{S}}}^{\mathcal{S}})\}$ and $\{\phi(c_{1}^{\mathcal{T}}),\cdots,\phi(c_{k_{\mathcal{T}}}^{\mathcal{T}})\}$. Then, the distance between the $i$-th class in the open-source datasets and the $j$-th class in the target dataset could be measured based on their cosine similarity:
${\rm cost}_{ij}=1-\frac{\phi(c_{i}^{\mathcal{S}})^{\top}\phi(c_{j}^{\mathcal{T}})}{\|\phi(c_{i}^{\mathcal{S}})\|\cdot\|\phi(c_{j}^{\mathcal{T}})\|}\;.$ (10)
Elements in ${\rm cost}\in\mathbb{R}_{+}^{k_{\mathcal{S}}\times k_{\mathcal{T}}}$ reveal how much the extracted text embeddings of these two classes are close to each other. In practical implementation, we utilize the exponential of each element in ${\rm cost}$ with the base of the natural constant $e$, which enlarges the difference of its values.
Values in the matrix ${\rm cost}$ indicate the extent of the cost required to reuse the element from a certain open-source class in a specific target class. In other words, if two classes are distant based on their textual semantics, it is more probable that reusing the statistics from that open-source class to help a target class is difficult. We take advantage of Optimal Transport (OT) to reuse more information between semantically similar classes:
@@
$\bm{\gamma}^{*}=\underset{\bm{\gamma}\in\mathbb{R}_{+}^{k_{\mathcal{S}}\times k_{\mathcal{T}}}}{\rm argmin}\sum_{i,j}\gamma_{i,j}\;{\rm cost}_{i,j}$ (11)
[…]
@@
The cost matrix ${\rm cost}$ quantifies the expense of moving elements between all class pairs, and $\bm{\gamma}^{*}\in\mathbb{R}^{k_{\mathcal{S}}\times k_{\mathcal{T}}}$ is the transport matrix. OT minimizes the cost indicated by the matrix ${\rm cost}$ and moves elements from one distribution $\bm{u}$ to another $\bm{v}$. In Swab, we define $\bm{u}$ and $\bm{v}$ as uniformly distributed vector $\bm{u}=\bm{1}/k_{\mathcal{S}}\in\mathbb{R}^{k_{\mathcal{S}}}$ and $\bm{v}=\bm{1}/k_{\mathcal{T}}\in\mathbb{R}^{k_{\mathcal{T}}}$. This indicates that we treat all classes as equally important. We may also incorporate prior knowledge of class importance to define $\bm{u}$ and $\bm{v}$.
@@
The solution $\bm{\gamma}^{*}$ of the OT problem in Equation 11 could be solved efficiently (Flamary et al., 2021), and $\bm{\gamma}^{*}$ acts as a bridge matrix between open-source classes and target classes. Usually, the smaller ${\rm cost}_{i,j}$ is, the larger the corresponding element $\bm{\gamma}^{*}_{i,j}$ obtained by OT, indicating statistics of the $i$-th open-source class may help more when we estimate the statistics of the $j$-th target class.
3.2 Bridge the Modality Gap and Capacity Gap
In Swab, we take advantage of the class relationship indicated by $\bm{\gamma}^{*}_{i,j}$ and bridge two kinds of gaps.
@@ 3.2 Bridge the Modality Gap and Capacity Gap
Bridge the Modality Gap. Given the $m$-th VLM $f_{m}$ in the model zoo, we want to estimate the modality gap $\bm{g}_{m,j}$ between the extracted image and text features for the $j$-th class in the target dataset $\mathcal{T}$ to bridge the modality gap. However, there are no images from the target dataset, so we can’t directly calculate the gap vectors using image-text pairs. Swab estimates the target dataset’s gap vectors based on the open-source datasets’ gap vectors with $\bm{\gamma}^{*}_{i,j}$.
Given the $k$-th open-source class $c_{k}^{\mathcal{S}}$, we collect the set of images $D_{\mathcal{S}_{k}}^{I}$ from this class
$D_{\mathcal{S}_{k}}^{I}=\left\{\left(\bm{x}_{i},y_{i}\right)|\left(\bm{x}_{i},y_{i}\right)\in D_{\mathcal{S}}^{I},\ y_{i}=c_{k}^{\mathcal{S}}\right\}.$ (12)
We denote $|D_{\mathcal{S}_{j}}^{I}|$ as the images number in $D_{\mathcal{S}_{j}}^{I}$. Then, the modality gap vector $\bm{g}_{m,k}^{\mathcal{S}}$ for $c_{k}^{\mathcal{S}}$ and $f_{m}$ could be obtained in a similar manner as Equation 9:
$\bm{g}_{m,k}^{\mathcal{S}}=\frac{1}{|D_{\mathcal{S}_{k}}^{I}|}\underset{(\bm{x}_{i},y_{i})\in D_{\mathcal{S}_{k}}^{I}}{\sum}\left(\frac{f_{m}^{I}\left(\bm{x}_{i}\right)}{\|f_{m}^{I}\left(\bm{x}_{i}\right)\|}-\frac{f_{m}^{T}\left(c_{k}^{\mathcal{S}}\right)}{\|f_{m}^{T}\left(c_{k}^{\mathcal{S}}\right)\|}\right)\;.$ (13)
$\bm{g}_{m,k}^{\mathcal{S}}$ is the average value of the difference between the normalized class name embedding and all normalized image embeddings from the class $c_{k}^{\mathcal{S}}$. Then, the gap vectors of all open-source classes $\{\bm{g}_{m,1}^{\mathcal{S}},\cdots,\bm{g}_{m,k_{\mathcal{S}}}^{\mathcal{S}}\}$ can be obtained given $f_{m}$. We use a matrix $\bm{G}_{m}^{\mathcal{S}}\in\mathbb{R}^{k_{\mathcal{S}}\times d_{m}}$ to represent those $k_{\mathcal{S}}$ gap vectors for the $m$-th VLM’s, and $d_{m}$ is the dimensionality of features extracted by $f_{m}$.
The gap vectors $\{\bm{g}_{m,1},\cdots,\bm{g}_{m,k_{\mathcal{T}}}\}$ for the target dataset could be estimated based on the $\bm{G}_{m}^{\mathcal{S}}$ and the class-wise relationship. If two classes are semantically similar, then we can reuse the gap vector from the similar class. We set the gap vector for the $j$-th target class $\bm{g}_{m,j}$ as a weighted sum of $\bm{G}_{m}^{\mathcal{S}}$, and the weight comes from $\bm{\gamma}^{*}_{i,j}$:
$\bm{g}_{m,j}=|C_{\mathcal{T}}|\;{\bm{G}^{\mathcal{S}}_{m}}^{\top}\bm{\gamma}^{*}_{:,j}.$ (14)
$\bm{\gamma}^{*}_{:,j}$ is the $j$-th column of $\bm{\gamma}^{*}$, whose elements are the relationship between the $j$-th target class to all open-source classes. We also use scaling factors $|C_{\mathcal{T}}|$ to ensure that for each target class, the sum of $\bm{\gamma}^{*}_{:,j}$ equals 1. This scale operation has also been used in previous work (Ye et al., 2018, 2021).
After that, we modify the essential step of ModelGPT in Equation 7, where the metrics over the generated auxiliary texts $\tilde{D}_{\mathcal{T}}$ are measured. We add the gap vector $\bm{g}_{m,j}$ to the embeddings of the auxiliary texts $\tilde{D}_{\mathcal{T}}^{j}$ from the $j$-th class in the target dataset:
$\tilde{\bm{t}}_{m,i}=f^{T}_{m}(t_{i})+\bm{g}_{m,j}\;,\;\;\forall t_{i}\in\tilde{D}_{\mathcal{T}}^{j}\;.$ (15)

### v3

@@ 2.3 Analysis of the Two Gaps in LOVM
There are two main challenges that limit the application of the aforementioned paradigms in LOVM. The first is the modality gap across different modalities’ features in VLM’s feature space, and the second is the capability gap between VLM’s overall performance and dataset-specific performance. Modality Gap. As described in Section 2.2, methods like H-Score, LogME, and ModelGPT utilize the ChatGPT generated auxiliary texts $\tilde{D}_{\mathcal{T}}$ as image proxies to calculate metrics that measure the zero-shot accuracy on the target dataset $\mathcal{T}$. In other words, the zero-shot classification ability across text and image modalities is estimated by the intra-modality classification ability. The latent assumption is that the generated texts and their corresponding images are closely aligned in VLM’s feature space. However, this assumption is difficult to meet [30], and instances’ features are more likely to cluster according to their modalities. In particular, we define the modality gap vector $\bm{g}$ between the features of an image-text pair $({\bm{x}}_{i},t_{i})$ as $\bm{g}_{m,i}\coloneqq f_{m}^{I}({\bm{x}}_{i})-f_{m}^{T}(t_{i})$. Values in the gap vector are generally not close to zero. We name this phenomenon as Modality Gap in LOVM, which makes the scores on $\tilde{D}_{\mathcal{T}}$ hard to reveal the true zero-shot image classification capability of a VLM on a given dataset.
@@
We conduct a validation experiment on ImageNet with 43 VLMs. We first generate 50 auxiliary texts per class as $\tilde{D}_{\mathcal{T}}$ and then calculate the predicted Top-1 accuracy via Equation 4. Next, we use test images to calculate the VLM’s true Top-1 accuracy. The consistency between the predicted Top-1 accuracy and true zero-shot image classification accuracy $p_{m,\mathcal{T}}$ is measured by the Kendall Rank Correlation ($\tau$, higher is better) and Mean Absolute Error (MAE, lower is better). It can be observed from the left part of Figure 2 that the predicted accuracy derived from auxiliary texts $\tilde{D}_{\mathcal{T}}$ does not closely match the true accuracy, indicating that these generated auxiliary texts in $\tilde{D}_{\mathcal{T}}$ are not effective proxies for images.
@@
To make the auxiliary texts act as better image proxies, one intuitive idea is to estimate the gap vector $\bm{g}$ for each image-text pair. Then we can add it to the feature $f_{m}^{T}(t_{i})$ of the text $t_{i}$ to eliminate the modality gap, which may lead to more accurate scores $s_{m,i}^{\mathcal{T}}$ in Equation 3. However, the gap vector cannot be calculated directly without the target dataset’s images. Furthermore, gap vectors for different classes are diverse, so using a shared vector across all datasets may not be a good choice.
Figure 2: Validation Experiments on the Modality Gap and Capability Gap. (a) Predicted VLMs’ zero-shot image classification accuracy based on generated text data vs. VLM’s true accuracy based on test images. Each point in the graph represents a model. From the result, we can find that the predicted accuracy poorly aligns with the true accuracy, indicating these text data are ineffective substitutes for image data. (b) We calculate the zero-shot image classification performance rankings of 43 VLMs across 23 datasets. We compute the average standard deviations and the mean value of differences between each VLM’s maximum and minimum ranking. The result shows the performance of a VLM varies greatly across different datasets.
@@
Capability Gap. To select one VLM from the model zoo given a target dataset, one direct approach is to select the VLM that performs the best on average across multiple datasets. For example, we may first estimate the VLM’s zero-shot classification ability on open-source datasets and then select the VLM with the highest performance. The key question is whether a VLM’s average ranking on the open-source datasets can reveal its true ranking on the target dataset. Our empirical analyses indicate that there exists a discrepancy between the VLM’s overall ranking and its ranking on a specific dataset. We name the discrepancy between the VLM’s average ability and its specific ability as the Capability Gap, which results from the fact that a VLM’s performance fluctuates significantly across various datasets.
@@
To verify the claim, we test 43 VLMs on 23 target datasets provided by [64] and obtain the rankings of each VLM across these datasets. Based on these ranking results, we calculate the average standard deviation and the mean value of the difference between each VLM’s maximum and minimum ranking. The experiment process is illustrated in the right part of Figure 2. We find that the mean difference between one VLM’s maximum and the minimum ranking is 34.86. Since the total number of VLMs is 43, such a difference demonstrates that the top-performing VLM in one dataset could likely be among the worst in another.
@@
One solution to bridge such a capability gap is to consider the VLMs’ ranking on a related dataset. In other words, the ranking of VLMs on datasets from open-source datasets collections that are relevant to the target task may provide more useful insights than a general performance ranking across all tasks. The main challenge is to figure out which open-source datasets are similar to the target dataset and transform the VLM’s ranking on these datasets to the target dataset.
@@
Summary. We emphasize two kinds of gaps in LOVM, i.e., the modality gap across features of different modalities generated by a VLM, and the capability gap between a VLM’s overall ranking and its ranking given a specific target dataset. Both two gaps pose obstacles to previous model selection methods, such as LogME and ModelGPT, and degrade their abilities in VLM selection. Moreover, those intuitive approaches to bridge the gaps still face challenges.
Figure 3: The workflow of Swab. Swab first constructs a transport matrix $\bm{\gamma}^{*}\in\mathbb{R}^{k_{\mathcal{S}}\times k_{\mathcal{T}}}$ using optimal transport, based on textual semantic similarity between classes in the open-source datasets $C_{\mathcal{S}}=\{c_{1}^{\mathcal{S}},\cdots,c_{k_{\mathcal{S}}}^{\mathcal{S}}\}$ and the target dataset’s classes $C_{\mathcal{T}}=\{c_{1}^{\mathcal{T}},\cdots,c_{k_{\mathcal{T}}}^{\mathcal{T}}\}$. Using this matrix, Swab estimates VLM $f_{m}$’s class-specific gap vectors $\{\bm{\hat{g}}_{m,1}^{\mathcal{T}},\cdots\}$ on the target dataset $\mathcal{T}$ from the gap vectors $\bm{G}_{m}^{\mathcal{S}}\in\mathbb{R}^{k_{\mathcal{S}}\times d}$ in the open-source datasets. These estimated gap vectors help modify text data to act as more effective substitutes for image data. The modified text data will then be input into the Ranker Model $f_{R}$, which predicts VLM’s performance $\hat{r}_{m}^{\mathcal{T},(1)}$ on the target dataset. Besides, Swab also uses the transport matrix $\bm{\gamma}^{*}$ to predict VLM’s performance ranking on the target dataset based on VLM’s class-specific rankings $\bm{r}_{m}^{\mathcal{S}}\in\mathbb{R}^{k_{\mathcal{S}}}$ on open-source datasets. Finally, Swab combines these two ranking predictions $\hat{r}_{m}^{\mathcal{T},(1)}$ and $\hat{r}_{m}^{\mathcal{T},(2)}$ to determine the VLM’s final ranking prediction.
@@ 3 VLM Selection with Gap Bridging
To mitigate the impact of both gaps on LOVM and integrate non-learning-based and learning-based LOVM methods, we propose VLM Selection With gAp Bridging (Swab). The key idea of Swab is to bridge modality and capability gaps by utilizing class-level statistics of VLMs from open-source datasets. By measuring the textual similarity between the target dataset’s class names and those in open-source datasets, we construct a bridge matrix. Based on it, we estimate the gap vectors between image and text modalities, which rectifies the text-derived scores in ModelGPT. In addition, we predict the VLM’s performance ranking for the target dataset based on the bridge matrix and VLM’s ranking on the open-source dataset. Both estimated statistics will be used to obtain a more accurate language-only VLM selection. The workflow of Swab is illustrated in Figure 3.
3.1 Construct the Bridge Matrix Using Optimal Transport
@@
Benefiting from the open-source datasets, some useful class-level statistics, such as modality gap vectors and zero-shot classification accuracy of a certain VLM, could be calculated, which can help the performance ranking estimation of a VLM on the target dataset. To better utilize these class-level statistics for predicting the corresponding statistics of a VLM on the target task, we introduce semantic relevance information between open-source datasets’ classes and target dataset’s classes into the statistics reusing process, which is automatically generated through Optimal Transport [7, 42].
@@
Recall that the sets of class names of the open-source datasets and the target dataset are $C_{\mathcal{S}}=\{c^{\mathcal{S}}_{i}\}^{k_{\mathcal{S}}}_{i=1}$ and $C_{\mathcal{T}}=\{c^{\mathcal{T}}_{i}\}^{k_{\mathcal{T}}}_{i=1}$, respectively. The semantic relevance between two classes could be measured by the textual similarity between their class names. In detail, we use a pre-trained text encoder $\phi$ (e.g., MPNet [46]), which extracts text features for these class names, i.e., $\{\phi(c_{1}^{\mathcal{S}}),\cdots,\phi(c_{k_{\mathcal{S}}}^{\mathcal{S}})\}$ and $\{\phi(c_{1}^{\mathcal{T}}),\cdots,\phi(c_{k_{\mathcal{T}}}^{\mathcal{T}})\}$. Then, we can calculate the cosine similarity $\frac{\phi(c_{i}^{\mathcal{S}})^{\top}\phi(c_{j}^{\mathcal{T}})}{\|\phi(c_{i}^{\mathcal{S}})\|\cdot\|\phi(c_{j}^{\mathcal{T}})\|}$ between the text feature of the $i$-th class in the open-source datasets $\phi(c_{i}^{\mathcal{S}})$ and that of the $j$-th class in the target dataset $\phi(c_{j}^{\mathcal{T}})$. The larger the cosine similarity, the more similar the two classes are. After that, we construct the cost matrix in optimal transport via ${\rm cost}_{ij}=1-\frac{\phi(c_{i}^{\mathcal{S}})^{\top}\phi(c_{j}^{\mathcal{T}})}{\|\phi(c_{i}^{\mathcal{S}})\|\cdot\|\phi(c_{j}^{\mathcal{T}})\|}$. In practice, we exponentiate each element in ${\rm cost}\in\mathbb{R}_{\geq 0}^{k_{\mathcal{S}}\times k_{\mathcal{T}}}$ using the base $e$ to amplify its differences. We then solve the optimal transport problem with the constructed cost matrix to get the transport matrix $\bm{\gamma^{*}}$. Since optimal transport aims to obtain a transport matrix that minimizes the transmission cost, the transport matrix $\bm{\gamma^{*}}$ will reuse more information between semantically similar classes:
@@
$\bm{\gamma}^{*}=\underset{\bm{\gamma}\in\mathbb{R}_{\geq 0}^{k_{\mathcal{S}}\times k_{\mathcal{T}}}}{\rm argmin}\sum_{i,j}\gamma_{i,j}\;{\rm cost}_{i,j}\;,\text{ s.t. }\bm{\gamma}\bm{1}=\bm{u};\ \bm{\gamma}^{T}\bm{1}=\bm{v};\ \gamma_{i,j}\geq 0.$ (5)
@@
The cost matrix quantifies the expense of moving elements between all class pairs, and $\bm{\gamma}^{*}\in\mathbb{R}^{k_{\mathcal{S}}\times k_{\mathcal{T}}}$ is the transport matrix. OT minimizes the cost indicated by the matrix ${\rm cost}$ and moves elements from one distribution $\bm{u}$ to another $\bm{v}$. In Swab, we define $\bm{u}$ and $\bm{v}$ as uniformly distributed vector $\bm{u}=\bm{1}/k_{\mathcal{S}}\in\mathbb{R}^{k_{\mathcal{S}}}$ and $\bm{v}=\bm{1}/k_{\mathcal{T}}\in\mathbb{R}^{k_{\mathcal{T}}}$. This indicates that we treat all classes as equally important. We may also incorporate prior knowledge of class importance to define $\bm{u}$ and $\bm{v}$.
@@
The solution $\bm{\gamma}^{*}$ of the OT problem in Equation 5 could be solved efficiently [13], and $\bm{\gamma}^{*}$ acts as a bridge matrix between open-source datasets’ classes and target dataset’s classes. Usually, the smaller ${\rm cost}_{i,j}$ is, the larger the corresponding element $\bm{\gamma}^{*}_{i,j}$ obtained by OT, indicating statistics of the $i$-th class of open-source datasets may help more when we estimate the statistics of the $j$-th target class.
3.2 Bridge the Modality Gap and Capability Gap
@@
Bridge the Modality Gap. Given the $m$-th VLM $f_{m}$ in the model zoo, we want to estimate the modality gap $\bm{g}_{m,j}^{\mathcal{T}}$ between the extracted image and text features for the $j$-th class in the target dataset $\mathcal{T}$ to bridge the modality gap. However, in the zero-shot scenario, we can’t get the target dataset’s images in advance, so we can’t directly calculate the gap vectors using image-text pairs. To solve this problem, Swab estimates the target dataset’s gap vectors based on the open-source datasets’ gap vectors with $\bm{\gamma}^{*}$. Given the $k$-th open-source class $c_{k}^{\mathcal{S}}$, we can get the set of images $D_{\mathcal{S}_{k}}^{I}=\left\{\left(\bm{x}_{i},y_{i}\right)|\left(\bm{x}_{i},y_{i}\right)\in D_{\mathcal{S}}^{I},\ y_{i}=c_{k}^{\mathcal{S}}\right\}$ from the open-source datasets. $|D_{\mathcal{S}_{k}}^{I}|$ is the number of images in $D_{\mathcal{S}_{k}}^{I}$. Then, the modality gap vector $\bm{g}_{m,k}^{\mathcal{S}}$ for class $c_{k}^{\mathcal{S}}$ and model $f_{m}$ can be calculated through $\bm{g}_{m,k}^{\mathcal{S}}=\frac{1}{|D_{\mathcal{S}_{k}}^{I}|}\underset{(\bm{x}_{i},y_{i})\in D_{\mathcal{S}_{k}}^{I}}{\sum}\left(\frac{f_{m}^{I}\left(\bm{x}_{i}\right)}{\|f_{m}^{I}\left(\bm{x}_{i}\right)\|}-\frac{f_{m}^{T}\left(c_{k}^{\mathcal{S}}\right)}{\|f_{m}^{T}\left(c_{k}^{\mathcal{S}}\right)\|}\right).$ $\bm{g}_{m,k}^{\mathcal{S}}$ is the average difference between the normalized class text prototype embedding and all normalized image embeddings from the class $c_{k}^{\mathcal{S}}$. In a similar manner, the gap vectors of all open-source classes $\{\bm{g}_{m,1}^{\mathcal{S}},\cdots,\bm{g}_{m,k_{\mathcal{S}}}^{\mathcal{S}}\}$ can be obtained given $f_{m}$. We use a matrix $\bm{G}_{m}^{\mathcal{S}}\in\mathbb{R}^{k_{\mathcal{S}}\times d_{m}}$ to represent those $k_{\mathcal{S}}$ gap vectors for the $m$-th VLM in the VLM Zoo, and $d_{m}$ is the dimensionality of features extracted by $f_{m}$.
The gap vectors $\{\bm{g}_{m,1}^{\mathcal{T}},\cdots,\bm{g}_{m,k_{\mathcal{T}}}^{\mathcal{T}}\}$ for the target dataset could be estimated based on the $\bm{G}_{m}^{\mathcal{S}}$ and the transport matrix $\bm{\gamma}^{*}$. If two classes are semantically similar, then we can reuse the gap vector from the similar class. We set the predicted gap vector for the $j$-th target class $\hat{\bm{g}}_{m,j}^{\mathcal{T}}$ as a weighted sum of $\bm{G}_{m}^{\mathcal{S}}$, and the weight comes from $\bm{\gamma}^{*}$, which is $\hat{\bm{g}}_{m,j}^{\mathcal{T}}=|C_{\mathcal{T}}|{(\bm{\gamma}^{*}_{:,j}})^{\top}\bm{G}^{\mathcal{S}}_{m}$ $\bm{\gamma}^{*}_{:,j}$ is the $j$-th column of $\bm{\gamma}^{*}$. We use scaling factors $|C_{\mathcal{T}}|$ to ensure that for each target class, the sum of $\bm{\gamma}^{*}_{:,j}$ equals 1. This scale operation has also been used in previous work [55, 56]. After that, we modify the step of ModelGPT in Equation 3, where the metrics over the generated auxiliary texts $\tilde{D}_{\mathcal{T}}$ are calculated. We add the gap vector $\hat{\bm{g}}_{m,j}^{\mathcal{T}}$ to the embeddings of the auxiliary texts $\tilde{D}_{\mathcal{T}}^{j}$ from the $j$-th class in the target dataset:
$\tilde{\bm{t}}_{m,i}=f^{T}_{m}(t_{i})+\hat{\bm{g}}_{m,j}^{\mathcal{T}}\;,\;\;\forall t_{i}\in\tilde{D}_{\mathcal{T}}^{j}\;.$ (6)